the equation x 5 – 2x 2 + 7 = 0 has (a) Atleast 2 imaginary roots (b) 2 negative real roots © 3 positive real roots (d) none of these ans:- (a) plzz give perfect explanation

the equation x5 – 2x2 + 7 = 0 has
(a) Atleast 2 imaginary roots
(b) 2 negative real roots
© 3 positive real roots
(d) none of these
ans:- (a)
plzz give perfect explanation

Grade:12

1 Answers

Vikas TU
7 years ago
Given equation x5 – 2x2 + 7 = 0 = f let
then
f’ = 5x^4 -4x = 0
x(5x^3 -4) = 0

plot the wavy curve and u will get 
for x > .983 slope increasing and 
for x
and    0
just after plotting the grapgh u will find that it has 2 real positive roots and one negative. and therfore remaing 2’s are imaginary.

hence (aP option.)